speed of wave = (frequency)(wavelength)

\begin{aligned}v_{s} & = fλ \\ c & = fλ \end{aligned}

f = \frac{1}{T}

T = \frac{1}{f}

(Δt)_{car} = \frac{1}{γ} (Δt)_{Fred}

T_{car} = \frac{1}{γ} T_{Fred}

f_{car} = γ f_{Fred}